# When can isotropy help adapt LLMs’ next word prediction to numerical domains?

### v3

@@ 5 Experiments
5.1 Qualitative Analysis
In this section, we analyze the time series forecasting by the baseline LLMs qualitatively. We focus on synthetically generated time series for a controlled analysis of different types of time series patterns which belong to $5$ different domains, such as linear, seasonality, trend, non-linear and stochastic. We are particularly interested in the isotropic measurement in the LLM’s last layer as it is related to the logits and probabilistic inference as explained in Section 2. So all isotropic measure provided in this section is based on the last layer of the baselines.
We begin by analyzing time series forecasting performance (i.e., NMSE) for different baselines and its relation with isotropy in Figure 6. For instance, in Figure 6 b, we have (NMSE = $0.0000066$ and cosine similarity = $|-0.00076|$) for seasonality-Dataset 1 and (NMSE = $0.00012$ and cosine similarity = $0.0047$) for seasonality-Dataset 2 for Chronos-T5. This shows stronger isotropy exists (i.e., inter-type cosine similarity value is close to 00) in Chronos-T5’s embedding space for seasonality-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists (i.e., inter-type cosine similarity value is far from 00) in Chronos-T5’s embedding space for seasonality-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to seasonality-Dataset 1. The NMSE and inter type cosine similarity may also vary for different language models and datasets. For example, in Figure 6 c, we can see that the the NMSE for trend-Dataset 1 is lower for PatchTST and Morai, while higher for Chronos-T5, Chronos-Bolt and Lag-Llma. In contrast, the NMSE for trend-Dataset 2 is lower for Chronos-T5, Chronos-Bolt, and Lag-Llma, while higher for PatchTST and Morai. A similar analysis can also be observed for other synthetic datasets and baselines in Figure 6 b, Figure 6 d, and Figure 6 e. This shows that any dataset from any particular domain may cause different forecasting performances for different baselines,
@@ 5.1 Qualitative Analysis
as it generates different hidden representations in contextual embedding spaces (see Appendix F for full visualization) for language models with different architectures and tokenization techniques.
Isotropy in different input context lengths. We first analyze the impacts of isotropy under varying input context lengths. In Figure 7, we compare the NMSE vs isotropy across two different input context lengths, $L=500$ and $L=100$, for different synthetic datasets. We use Choronos-T5 as an example model for this experiment. As can be seen from the figure, the isotropy values vary across different input context lengths and datasets. For instance, in seasonality-Dataset 1, we have (NMSE$=0.0000066$, cosine similarity$=|-0.00076|$) and (NMSE$=0.0793$, cosine similarity$=0.0011$) for $L=500$ and $L=100$, respectively. The decrease in isotropy significantly increases the NMSE for the input context length $L=100$. In contrast, in linear-Dataset 2, we have (NMSE$=0.000025$, cosine similarity$=0.2474$) and (NMSE$=0.000009$, cosine similarity$=0.0644$) for $L=500$ and $L=100$, respectively. In this scenario, the isotropy increases for the input context length $L=100$, which causes the decrease in NMSE for chornos-T5. In Figure 11 in Appendix F, we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=500$, for non-linear-Dataset 1, which generates different isotropic measures for different input context lengths. In practice, the input context length is often selected randomly or through trial and error, which may cause higher forecasting errors for different datasets. Isotropy analysis enables us to understand how varying input context lengths influence the hidden representations of the language model. This insight helps guide improvements in forecasting performance by examining the isotropic properties of the contextual embedding space.
@@
Isotropy in varying noise levels in datasets. Next, we focus on the second scenario to see the impact of noisy datasets on LLM’s performance. Again, we use the Chornos-T5 as an example language model. In Figure 8, we compare the NMSE vs isotropy across two different noise levels, one without noise, and the other with Gaussian noise with $\sigma=0.05$ standard deviation. From Figure 8, we can see consistently lower isotropy (i.e., inter-type cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise. For instance, in trend-Dataset 2, we have (NMSE$=0.000024$, cosine similarity$=|-0.00022|$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively. The decrease in isotropy significantly increases the NMSE for the noisy dataset. In practice, many real-world numerical domains—such as those in nature and energy—exhibit noisy and dynamic behavior. In these environments, it is often infeasible to measure noise in real time or to pre-process the input time series for improved performance. However, the isotropy in LLMs hidden representations can be readily measured, and thus, can be leveraged to enhance forecasting performance by identifying and mitigating the effects of noisy inputs in contextual embedding space.
@@ 5.2 Quantitative Analysis
In this section, we present our main results on $12$ real datasets which belong to $6$ different numerical domains such as energy, nature, finance, healthcare, retail and transportation. As our qualitative analysis in Section 5.1, we select two different datasets from each numerical domain and the isotropy measure from LLM’s last layer to show the impact of isotropy on NMSE performance for different language models.
In Figure 9, we analyze the time series forecasting performance of different baselines and its relation with isotropy for different real datasets. For instance, in Figure 9 e, we have (NMSE = $0.0061$ and cosine similarity = $0.0020$) for retail-Dataset 1 and (NMSE = $0.1255$ and cosine similarity = $0.1931$) for retail-Dataset 2 for Morai. This indicates the existence of stronger isotropy in Morai’s embedding space for retail-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists in Morai’s embedding space for retail-Dataset 2, and thus causes a lack of structure in its hidden representations which leads to bad downstream task performance as compared to retail-Dataset 1. The NMSE and inter type cosine similarity may also vary for different real datasets and language models. For example, in Figure 9 a, we can see that the NMSE for energy-Dataset 1 is lower for Chronos-T5, Chronos-Bolt, PatchTST and Morai, while higher Lag-Llma. In contrast, the NMSE for energy-Dataset 2 is lower for Morai while higher for other baselines. A similar analysis can also be observed for other synthetic datasets and baselines in Figure 9 b and Figure 9 c. This again shows that datasets from the same numerical domain can cause varying forecasting performance across different baselines, as they generate distinct hidden representations in contextual embedding spaces depending on the language model architecture and tokenization strategy.
@@
Finally, in Figure 10, we compare the NMSE vs isotropy for varying input context lengths to observe its impact on the real datasets. We select Lag-Llma as our example model. We compare the results for two different input context lengths: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$. As can be seen from the figure, the inter-type cosine similarity values become far from 00, i.e., from $0.0097$ to $0.0220$ for energy-Dataset 1 and from $0.0026$ to $0.0103$ for transport-Dataset 1, which in turn decreases the NMSE performances. On the other hand, the inter-type cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0112$ for nature-Dataset 2 and from $0.2014$ to $0.0133$ for finance-Dataset 1, which in turn improves the NMSE performances. Thus, the variation in the recommended input context length may not only decrease the NMSE performances, but can also increase for some datasets.
@@ 6 Conclusion and Limitations
Isotropy in embeddings as studied here can serve as a foundation for future research on the deeper understanding of LLMs and their applications in various domains. Beyond isotropy, there could be other methods to approximate the partition function with a constant and make the logits useful for the numeric downstream tasks. Moreover, our isotropy study only ensured the existence of structure in the LLM hidden representations and provides a performance guarantee when the structure is preserved by isotropy. Improving the numeric downstream task performance when structure is not preserved in the LLM representations is a topic of future work.
Appendix A Proof of Theorem 3.1
Theorem A.1.
Let the logits of the ground-truth model be bounded. Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$. In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numeric downstream tasks.
@@
Appendix B Proof of Lemma 4.1
Lemma B.1.
@@
The theorem below shows that $\mathbb{\Lambda}$ minimizing the objective $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ contains the largest $m$ eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of input time series token embeddings where $m$ is the rank of $\mathbb{\Lambda}$.
@@
To proceed, we consider the objective in the scenario where $\rho$ is small. In this case, we can approximate the attention weights by $p_{i,j}\approx\frac{1}{|\mathcal{V}|}+\frac{1}{|\mathcal{V}|}\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{j}$. Now, we define the average of embedding as $\bar{\mathbb{\psi}}=\mathbb{\Psi}^{\top}1/|\mathcal{V}|$. It then follows that $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}=\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\bar{\mathbb{\psi}}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$. Assuming all input time series patterns are zero-centered, i.e., $\bar{\mathbb{\psi}}=0$, we have $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}=\text{tr}\left((I-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda})^{2}\mathbb{\Psi}^{\top}\mathbb{\Psi}\right)$. Theorem 4.1 establishes that the optimal $\mathbb{\Lambda}$ that minimizes $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ is spanned by the top $m$ eigenvectors of $\mathbb{\Psi}^{\top}\mathbb{\Psi}$, where $m$ equals the rank of $\mathbb{\Lambda}$.
Appendix C Proof of Theorem 4.1
Theorem C.1.

### v4

@@ 5.1 Qualitative Analysis
We begin by analyzing time series forecasting performance (i.e., NMSE) for different baselines and its relation with isotropy in Figure 3. For instance, in Figure 3 b, we have (NMSE = $0.0000066$ and cosine similarity = $|-0.00076|$) for seasonality-Dataset 1 and (NMSE = $0.00012$ and cosine similarity = $0.0047$) for seasonality-Dataset 2 for Chronos-T5. This shows that stronger isotropy exists (i.e., inter-token cosine similarity value is close to 00) in Chronos-T5’s embedding space for seasonality-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists (i.e., inter-token cosine similarity value is far from 00) in Chronos-T5’s embedding space for seasonality-Dataset 2, which, in turn, causes a lack of structure in its hidden representations, thereby leading to bad forecasting performance as compared to seasonality-Dataset 1. The NMSE and inter-type cosine similarity can also vary across different language models and datasets. For example, in Figure 3c, the NMSE for trend-Dataset 1 is lower for PatchTST and Moirai, but higher for Chronos-T5, Chronos-Bolt, and Lag-Llama, compared to their respective NMSE on trend-Dataset 2. Conversely, for trend-Dataset 2, the NMSE is lower for Chronos-T5, Chronos-Bolt, and Lag-Llama, but higher for PatchTST and Moirai, compared to their respective NMSE on trend-Dataset 1. A similar analysis can also be observed for other synthetic datasets and baselines in Figures 3 b, 3 d, and 3 e.
This shows that any dataset from any particular domain may cause different forecasting performances for different baselines, as it generates different hidden representations (see Appendix F for full visualization) in contextual embedding spaces, and hence, different isotropy measures.
Isotropy in different input context lengths. We first analyze the effect of isotropy under varying input context lengths. We begin with an illustration in Figure 4 where we show how the hidden representations of Chronos-T5 vary for two different input context lengths, such as $L=500$ and $L=100$, for non-linear-Dataset 1, which generates different isotropic measures for different input context lengths.
@@
In Figure 5, we compare the NMSE vs isotropy across two different input context lengths, $L=500$ and $L=100$, for different synthetic datasets. We use Choronos-T5 as an example model for this experiment. As can be seen from the figure, the isotropy values vary across different input context lengths and datasets. For instance, in seasonality-Dataset 1, we have (NMSE$=0.0000066$, cosine similarity$=|-0.00076|$) and (NMSE$=0.0793$, cosine similarity$=0.0011$) for $L=500$ and $L=100$, respectively. The decrease in isotropy significantly increases the NMSE for the input context length $L=100$. In contrast, in linear-Dataset 2, we have (NMSE$=0.000025$, cosine similarity$=0.2474$) and (NMSE$=0.000009$, cosine similarity$=0.0644$) for $L=500$ and $L=100$, respectively. In this scenario, the isotropy increases for the input context length $L=100$, which causes the decrease in NMSE for chornos-T5. In practice, the input context length is often selected randomly or through trial and error, which may cause higher forecasting errors for different datasets. Isotropy analysis enables us to understand how varying input context lengths influence the hidden representations of the language model. This insight helps guide improvements in forecasting performance by examining the isotropic properties of the contextual embedding space.
@@
Isotropy in varying noise levels in datasets. Next, we focus on the second scenario to see the impact of noisy datasets on LLM’s performance. Again, we use the Chornos-T5 as an example language model. Figure 6 compares the NMSE vs isotropy across two different cases, one without noise, and the other with Gaussian noise with a standard deviation $\sigma=0.05$ standard deviation. From Figure 6, we can see consistently lower isotropy (i.e., inter-token cosine similarity far from 00) for all noisy synthetic datasets as compared to the datasets without noise. For instance, in trend-Dataset 2, we have (NMSE$=0.000024$, cosine similarity$=|-0.00022|$) and (NMSE$=0.0012$, cosine similarity$=0.0040$) for $\sigma=0$ and $\sigma=0.05$, respectively. The decrease in isotropy significantly increases the NMSE for the noisy dataset. In practice, many real-world numerical domains—such as those in nature and energy—exhibit noisy and dynamic behavior.
In these environments, it is often infeasible to measure noise in real time or to pre-process the input time series for improved performance. However, the isotropy in the hidden representations of LLMs can be readily measured, and thus, can be leveraged to enhance forecasting performance by identifying and mitigating the effects of noisy inputs in contextual embedding space.
@@ 5.2 Quantitative Analysis
Next, we present our main results on $12$ real datasets which belong to $6$ different numerical domains including energy, nature, finance, healthcare, retail, and transportation. As our qualitative analysis in Section 5.1, we select two different datasets from each numerical domain and the isotropy measure from LLM’s last layer to show the impact of isotropy on NMSE performance for different language models.
In Figure 7, we analyze the time series forecasting performance of different baselines and its relation with isotropy for different real datasets. For instance, in Figure 7 e, we have (NMSE = $0.0061$ and cosine similarity = $0.0020$) for retail-Dataset 1 and (NMSE = $0.1255$ and cosine similarity = $0.1931$) for retail-Dataset 2 for Moirai. This indicates the existence of stronger isotropy in Moirai’s embedding space for retail-Dataset 1 which preserves the structure in its hidden representations and causes good downstream task performance. On the other hand, a weaker isotropy exists in Moirai’s embedding space for retail-Dataset 2, which yields a lack of structure in its hidden representations and, consequently, bad downstream task performance as compared to retail-Dataset 1. The NMSE and inter-type cosine similarity can vary across different real datasets and language models. For example, in Figure 7a, the NMSE for energy-Dataset 1 is lower for Chronos-T5, Chronos-Bolt, PatchTST, and Moirai, but higher for Lag-Llama, compared to their respective NMSE on energy-Dataset 2. Conversely, the NMSE for energy-Dataset 2 is
@@
lower for Moirai but higher for the other baselines, compared to their respective NMSE on energy-Dataset 1. A similar analysis can also be observed for other synthetic datasets and baselines in Figure 7 c and 7 e. This again shows that datasets from the same numerical domain can cause varying forecasting performance across different baselines, as they generate distinct hidden representations in contextual embedding spaces, and hence, different isotropy measures, depending on the language model architecture and tokenization strategy.
Finally, in Figure 8, we compare the NMSE vs isotropy for varying input context lengths to observe its impact on the real datasets. We select Lag-Llma as our example model. We compare the results for two different input context lengths: 1) the recommended input context length $L=144$ and the reduced input context length $L=96$. As can be seen from the figure, the inter-token cosine similarity values become far from 00, i.e., from $0.0097$ to $0.0220$ for energy-Dataset 1 and from $0.0026$ to $0.0103$ for transport-Dataset 1, which in turn decreases the NMSE performances. On the other hand, the inter-token cosine similarity values become close to 00, i.e., from $0.1091$ to $0.0112$ for nature-Dataset 2 and from $0.2014$ to $0.0133$ for finance-Dataset 1, which in turn improves the NMSE performances. Thus, the variation in the recommended input context length may not only decrease the NMSE performances, but can also increase for some datasets.
@@ 6 Conclusion and Limitations
In this work, we introduced a novel approach to investigate the role of isotropy in LLM hidden representations for numerical downstream tasks. By deriving an upper bound for the Jacobian matrix which collects all first-order partial derivatives of self-attention with respect to the input pattern, we showed that the self-attention mechanism implicitly aligns with the dominant eigenvectors of the input correlation structure and induces isotropy in the contextual embedding space. The existence of isotropy in the contextual embedding space was found to stabilize the partition function and enable better generalization in numerical downstream tasks across different models and datasets. Our empirical analysis across $10$ synthetic and $12$ real numerical datasets, and $5$ different language models further validated the consistent relationship between isotropy and forecasting performance, highlighting isotropy as a reliable indicator of structured representation learning. These insights open up a new interpretability frontier for LLMs in numerical domains.
While isotropy offers a principled way to preserve useful structure, there may be alternative approaches to approximating the partition function and guiding numerical reasoning. Moreover, developing mechanisms to recover or enhance structure when isotropy is weak remains an important avenue for future work. Ultimately, we believe that incorporating structural insights like isotropy into the LLM design pipeline can significantly improve their reliability and adaptability to numerical domains.
Appendix A Proof of Theorem 1
Theorem 1. Let the logits of the ground-truth model be bounded. Then for any $f^{*}(k,l)$, there exists a set of functions $\{\hat{z}_{i}(k,l)\}_{i=1}^{|\mathcal{V}|}$ such that for all $k$ and $T_{l+1}$, the predictive distribution of the student model $\hat{p}\,(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ matches that of ground-truth model $p^{*}(k_{T_{l+1}}\,|\,\mathbb{k}_{1:T_{l}})$ and $\hat{f}(k,l)=0$. In other words, there exists a student model with the same pre-training loss as the ground-truth model, but its logits are ineffective for the numerical downstream tasks.
@@ Appendix A Proof of Theorem 1
Appendix B Proof of Lemma 1
Lemma 1. Consider the Jacobian matrix $\mathbb{J}=\left[\frac{\partial g_{i}(\mathbb{\Psi})}{\partial\mathbb{\psi}_{j}}\right]_{i,j=1}^{|\mathcal{V}|}$, which represents the gradient of the self-attention mapping $G(\mathbb{\Psi})$ with respect to the input time series token embeddings. Then the spectral norm of $\mathbb{J}$ satisfies $\|\mathbb{J}\|_{2}\leq|\mathbb{\Lambda}|_{2}\sum_{i=1}^{|\mathcal{V}|}\left(p_{i,i}+\frac{1}{2}\right)\left|\mathbb{\psi}_{i}-\sum_{j=1}^{|\mathcal{V}|}p_{i,j}\mathbb{\psi}_{j}\right|^{2}+\Delta$, where the residual term $\Delta$ is given by $\Delta=|\mathbb{\Lambda}|_{2}\sum_{i\neq j}^{|\mathcal{V}|}p_{i,j}\left|\mathbb{\psi}_{j}-\sum_{q=1}^{|\mathcal{V}|}p_{i,q}\mathbb{\psi}_{q}\right|^{2}+\frac{|\mathbb{\Lambda}|_{2}}{2}\sum_{j=1}^{|\mathcal{V}|}|\mathbb{\psi}_{j}|^{2}$, and the attention weights $p_{i,j}$ are defined as $p_{i,j}=\frac{\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{j})}{\sum_{k=1}^{|\mathcal{V}|}\exp(\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{k})}$.
@@ Appendix B Proof of Lemma 1
Theorem 2 shows that $\mathbb{\Lambda}$ minimizing the objective $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ contains the largest $m$ eigenvectors of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ of input time series token embeddings where $m$ is the rank of $\mathbb{\Lambda}$.
@@
To proceed, we consider the objective in the scenario where $\rho$ is small. In this case, we can approximate the attention weights by $p_{i,j}\approx\frac{1}{|\mathcal{V}|}+\frac{1}{|\mathcal{V}|}\mathbb{\psi}_{i}^{\top}\mathbb{\Lambda}\mathbb{\psi}_{j}$. Now, we define the average of embedding as $\bar{\mathbb{\psi}}=\mathbb{\Psi}^{\top}1/|\mathcal{V}|$. It then follows that $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}p_{i,:}|^{2}=\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\bar{\mathbb{\psi}}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$. Assuming all input time series patterns are zero-centered, i.e., $\bar{\mathbb{\psi}}=0$, we have $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}=\text{tr}\left((I-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda})^{2}\mathbb{\Psi}^{\top}\mathbb{\Psi}\right)$. Theorem 2 establishes that the optimal $\mathbb{\Lambda}$ that minimizes $\sum_{i=1}^{|\mathcal{V}|}|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}|^{2}$ is spanned by the top $m$ eigenvectors of $\mathbb{\Psi}^{\top}\mathbb{\Psi}$, where $m$ equals the rank of $\mathbb{\Lambda}$.
Appendix C Proof of Theorem 2
Theorem 2. Let the eigenvalues of the correlation matrix $\mathbb{\Psi}^{\top}\mathbb{\Psi}$ be ordered as $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{D}$, and let $\gamma_{i}\in\mathbb{R}^{D}$ for $i=1,\ldots,D$ denote their associated eigenvectors. Then, the matrix $\mathbb{\Lambda}^{*}$ that minimizes the quantity $\sum_{i=1}^{|\mathcal{V}|}\left|\mathbb{\psi}_{i}-\mathbb{\Psi}^{\top}\mathbb{\Psi}\mathbb{\Lambda}\mathbb{\psi}_{i}\right|^{2}$ has the optimal form $\mathbb{\Lambda}=\sum_{i=1}^{m}\frac{1}{\lambda_{i}}\gamma_{i}\gamma_{i}^{\top}.$